OpenStudy (anonymous):

interval of convergence of... (-1)^n*(x+2)^n / n^2, n goes from 0 to infinity. I know you use the ratio test but how do you get rid of the x before you take the limit??

OpenStudy (anonymous):

i need help im really bad at math and i still cant figure out slope intercept form can you help me on this problem: i have to write an equation in slope intercept form that passes through the given points. 2,-2 , (5,7)

OpenStudy (anonymous):

You don't. You take the limit and treat the x as a constant.

OpenStudy (anonymous):

Then you get (whatever after limit)<|1| Then you use that to solve for the radius.

OpenStudy (anonymous):

Errr...|whatever after limit|<1** (what it should read)

OpenStudy (anonymous):

What did you get for the ration n+1/n? (before you take the limit)

OpenStudy (anonymous):

lim of -n^2/(n+1)^2 *(x+2)

OpenStudy (anonymous):

Yes. Or written as: \[\lim_{n \rightarrow \infty} |(\frac{n}{n+1})^2 (x+2)|<1\]. Since you take the abs value you can ignore the -.So taking the limit you can divide the ()^2 by n. \[\lim_{n \rightarrow \infty} |(\frac{1}{1+\frac{1}{n}})^2 (x+2)|<1\]. That limit gives you |x+2|<1 Or: -1<x+2<1 -3<x<-1.

OpenStudy (anonymous):

Tell me if you need elaboration :P

OpenStudy (anonymous):

got it! thanks!

OpenStudy (anonymous):

No problem :D

OpenStudy (anonymous):

Then, you need to plug in the endpoints in for x. See if the series converges, if it DOES then it has a closed interval for that x value. If the series DIVERGES, then its open. For example: if -3 made the series converge, and the -1 made it diverge. You would have: [-3,-1) (I'm not sure which one either does xP)

OpenStudy (anonymous):

oo rite..always forget that step! thanks again

OpenStudy (anonymous):

Haha, no problem :P

OpenStudy (anonymous):

wat does the radius of convergence mean?

OpenStudy (anonymous):

So you have -3 to -1 they are 2 units apart right? In other words -1-(-3)=2 The radius is half of the interval, or 1.